### v1

# Neural Differential Algebraic Equations

Abstract
Differential-Algebraic Equations (DAEs) describe the temporal evolution of systems that obey both differential and algebraic constraints. Of particular interest are systems that contain implicit relationships between their components, such as conservation relationships. Here, we present Neural Differential-Algebraic Equations (NDAEs) suitable for data-driven modeling of DAEs. This methodology is built upon the concept of the Universal Differential Equation; that is, a model constructed as a system of Neural Ordinary Differential Equations informed by theory from particular science domains. In this work, we show that the proposed NDAEs abstraction is suitable for relevant system-theoretic data-driven modeling tasks. Presented examples include (i) the inverse problem of tank-manifold dynamics and (ii) discrepancy modeling of a network of pumps, tanks, and pipes. Our experiments demonstrate the proposed method’s robustness to noise and extrapolation ability to (i) learn the behaviors of the system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the system.
I INTRODUCTION
Modeling and simulation of systems via Differential-Algebraic Equations (DAEs) can be difficult, both in terms of formulation as well as numerical implementation. Common issues include (i) over- or under-constraining states and (ii) incorrect model specification, either of which can lead to a prohibitively challenging modeling task [1]. While many such algebraic relationships may be known (e.g., mass is conserved through a pipe), the relationships between system states and these algebraic constraints are often difficult to reconcile in a modeling framework (see, for example, index of a DAE and choosing appropriate initial conditions [1]). Incorporating these algebraic relationships into well-established data-driven modeling paradigms is also non-trivial. Here, measurement noise, partial observations, and missing or incorrect physics all confound the ability to effectively build models that obey algebraic constraints.
Furthermore, it is generally rare to construct a DAE-based model of an engineering system from first principles. More common is to adopt simplified models that can capture improperly specified or missing physics at a given fidelity. A common approach is to employ a closure model [2], i.e., a surrogate for the full-fidelity physics inserted into a lower-fidelity model. Closure models have enabled major advancements in scientific computing but still traditionally rely upon rigorous reduction from first principles physics to derive required simplified relationships [3].
To address this issue, we present a framework for the data-driven modeling of DAEs with a novel neural timestepper scheme. Our framework leverages themes from operator splitting methods [4] and Physics-Informed Neural Network (PINN)-style [5] optimization problems to accommodate algebraic relationships. We demonstrate the utility of these methods on two closure modeling tasks: (i) a tank-manifold problem, where one needs to ‘invert’ the model to provide an estimate of a property of a model component, and (ii) a tank-pipe-pump network for learning unknown nonlinear interactions. Our specific contributions are:
  • Taking inspiration from the well-established field of operator splitting, we extend neural timesteppers to appropriately integrate DAEs.
  • We accomplish this via an algebra surrogate layer that modifies algebraic states of a system, while an ODE Solver integrates differential states.
  • We demonstrate the utility of the method in applications to canonical closure and inverse modeling of constrained networked system dynamics.
The paper is organized as follows. In Section II, we give a brief overview of recent methods proposed for data-driven modeling of agent-based systems. Section III details our proposed methodology and its implementation. Example results are provided in Section IV, followed by a discussion in Section V.
II BACKGROUND
II-A Differential-Algebraic Equations
Differential-Algebraic Equations (DAEs) are systems comprising differential and algebraic relationships that describe the evolution of a system’s states. Specifically, we are concerned with semi-explicit DAEs written as:
$\begin{split}\frac{dx}{dt}&=f\left(x,y,u\right)\\ 0&=g\left(x,y,u\right),\end{split}$ (1)
where $x\in\mathbb{R}^{n_{x}}$ are the differential states of a system that evolve according to the vector field defined by $f:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{x}}$, $y\in\mathbb{R}^{n_{y}}$ are the algebraic states, and $u\in\mathbb{R}^{n_{u}}$ are exogenous inputs. The algebraic states $y$ evolve such that the algebraic relationships, defined by $g(x,y,u)=0$, are satisfied for all time $t$.
A standard DAE solution approach is to transform the system into a set of ODEs to be integrated with standard numerical solvers. This process is termed index reduction [1]. The index of a DAE is a distance measure between a DAE and its associated ODE: an index-1 DAE requires one differentiation of algebraic relationships to yield a consistent set of ODEs, index-2 requires two differentiations, etc. After differentiation and substitution into other system relationships, the governing equations can be written as a set of first-order ODEs.
II-B Neural Differential Equations
A Neural Ordinary Differential Equation (NODE) approximates the flow map of a dynamical system with a neural network architecture. Consider a dynamical system with states $x\in\mathbb{R}^{n_{x}}$. A NODE approximates the temporal dynamics of $x$ according to the ODE:
$\frac{dx}{dt}=f\left(x;\theta\right),$ (2)
where the parametric flow $f:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{x}}$ maps the states to their time derivatives with tunable parameters $\theta$. The solution to Eq. (2) is the initial value problem:
[…]
where the superscript contained in parentheses denotes indexing by time, not exponentiation. Practically, this integration is performed with a standard numerical ODE solver, which we denote as:
$x^{(t+\Delta t)}=\operatorname{ODESolve}\left(f,x^{(t)};\theta\right),$ (4)
where the limits of temporal integration are understood to be $[t,t+\Delta t]$. In general, the NODE concept extends $f$ to any differentiable mapping by implementing Eq. (4) in a differentiable manner [6]. This is achieved through either (i) backpropagating residuals through the elementary functions of the explicit time-stepping integrator (reverse-mode automatic differentiation) [7, 8], (ii) an adjoint sensitivity definition and solution [9], or (iii) forward-mode automatic differentiation [10].
A fundamental assumption of a NODE model is that a direct relationship exists between each system state and its temporal derivative. This is a reasonable assumption for many cases, but poses challenges for partially-observed systems [11] and application to DAEs. Index reduction is a reliable technique for recasting a DAE into a set of ODEs for standard numerical integration [12], but this is incompatible with the present data science problem. When governing equations are unknown or incomplete, the equivalent set of ODEs and required state augmentations are also unknown.
II-C Fractional Step Methods
Our methodology is inspired by fractional step methods for evolving complex multiphysics problems. In fractional step methods (also termed ‘Operator Splitting’), the evolution of a system is split into simpler, independent sub-tasks. The underlying ODE or PDE is then solved explicitly as the sequential solution of these sub-tasks, albeit at a decreased level of accuracy (i.e., splitting error [13]). There are several well-established mathematical techniques for decomposing the solution of an ODE into sub-tasks. The simplest scheme, Lie-Trotter operator splitting [4], assumes a differential operator can be decomposed into two sub-tasks. For the ODE of Eq. 2, assume $f$ can be decomposed into operators $A$ and $B$ such that:
$\frac{dx}{dt}=A(x)+B(x)$ (5)
for the interval $t\in[t,t+\Delta t]$. The Lie-Trotter splitting scheme defines the sub-tasks as:
$\hat{x}^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(A(x),x^{(t)}\right)$ (6)
$x^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(B(x),\hat{x}^{(t+\Delta t)}\right)$ (7)
The first sub-task solves the ODE defined by the operator $A(x)$ on the time interval $[t,t+\Delta t]$. The second sub-task solves the ODE defined by the operator $B(x)$ on the same time interval, but with the initial conditions defined by the solution of the first sub-task. Many variations of this scheme exist for different applications and orders of accuracy. For example, Strang splitting [14] is a second-order accurate scheme that uses half-steps (i.e., steps of $\frac{\Delta t}{2}$) in the above scheme.
II-D Related Work
II-D1 Machine Learning for Dynamical Systems
Methods for learning dynamics from data exist on a continuum from ‘black-box’ models to ’white-box’ models [15]. The ‘black-box’ modeling task is performed without regard to the underlying system or domain knowledge and includes Recurrent Neural Networks (RNNs) and their variants (e.g., LSTMs)  [16], neural state space models [17], neural timesteppsers  [8], sparse symbolic regression [18], and Neural ODEs  [9]. Often, such models extract salient features in the training data to infer future states. However, they are at risk of poor generalization (overfitting) and lack of interpretability, except for symbolic regression techniques [18]. ‘White-box’ models use domain knowledge to constrain models to match expected dynamics by construction and often lead to parameter-tuning problems [19]. In ‘gray-box’ models, recently referred to as physics-informed machine learning [20], prior knowledge is incorporated in a manner to constrain a ‘black-box model’. Neural ODEs constructed with domain knowledge, termed Universal Differential Equations  [6, 21], are an attractive modeling platform for these problems because of their flexibility in encoding physics and other structural priors.
II-D2 Data-driven Modeling of DAEs
If the DAE is well-characterized, the parameter estimation, inverse design, and control tasks are possible through an adjoint sensitivity analysis  [22, 23, 24]. This strategy is generally successful if (i) the index of the underlying DAE is low and (ii) all constituent physics are known.
Recent works have begun to address some common issues regarding data-driven DAE-based models. Moya and Lin  [25] use explicit Runge-Kutta time-stepping to combine semi-explicit DAEs into machine learning frameworks, which they demonstrate by modeling power networks. Xiao et al.  [26] construct neural ODE and DAE modules to forecast a regional power network. Huang et al.  [27] adopt a simplified approach to the construction of a surrogate system for sequence-to-sequence mappings for DAEs. These studies share features of expressing the learning task in terms of a semi-explicit DAE with varying degrees of parameterization and integration of domain knowledge. In addition to these studies, Zhong et al.  [28] addressed contact modeling in multi-body dynamics, the authors in the work handled a constraint update architecture with similar explicit time-stepping as in Moya  [25] and Xiao  [26].
While the specific treatment of each of these domain-focused problems varies, each leverages a form of a multi-step integration technique that separates the evolution of differential and algebraic states. Here, we recognize that the commonality of these works holds the potential for a more powerful generalization of standard neural integration. Specifically, we recast these ideas under the umbrella of fractional-step methods, where integration is split into simpler, sequential sub-tasks. In effect, we propose a new class of neural integrators that are well-suited for a wide variety of modeling tasks, including DAEs, boundary value problems, and other multi-physics systems.
III METHODS
Consider the DAE system of Eq. 1. Similar to the approach of a Lie-Trotter fractional step scheme, we split the evolution of a system’s differential and algebraic states into sequential sub-tasks. The first sub-task updates the algebraic states using a trainable deep neural network (DNN). The second sub-task takes as input the updated algebraic states and performs a single-step integration. These two sequential sub-tasks define one complete forward pass; all system states are updated exactly once. More formally, consider a differentiable ODE layer, denoted as $\operatorname{ODESolve}$. Using a fractional step method as our solution technique for Eq. 1, we formulate a Neural DAE (NDAE) as:
$\begin{split}y^{(t+\Delta t)}&=h\left(x^{(t)},y^{(t)},u^{(t)};\theta_{h}\right)\\ x^{(t+\Delta t)}&=\operatorname{ODESolve}\left(f,\{x^{(t)},y^{(t+\Delta t)},u^{(t)}\};\theta_{f}\right),\end{split}$ (8)
where $h:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{y}}$ pairs with $g$ as an explicit algebra solver or trainable neural surrogate thereof and the temporal integration bounds are assumed to be $[t,t+\Delta t]$. We denote this complete set of operations as $\operatorname{DAESolve}$, with one complete time step captured as:
$\begin{split}\{x^{(t+\Delta t)},y^{(t+\Delta t)}\}&=\\ \operatorname{DAESolve}&\left(f,h,\{x^{(t)},y^{(t)},u^{(t)}\};\Theta\right),\end{split}$ (9)
where $\Theta=\{\theta_{h},\theta_{f}\}$ is the set of all model parameters.
Note that in this formulation, the updates to the algebraic and differential states are sequential; that is, the $\operatorname{ODESolve}$ step is dependent upon the update to the algebraic states. The sub-tasks do not commute. This is an essential assumption that enables explicit updates for both algebraic states and differential states. Figure 1 depicts the handling of differential and algebraic states through the Neural DAE. The module generates an internal state $y^{(t+\Delta t)}$ that is processed by the subsequent $\operatorname{ODESolve}$ as an exogenous input, similar to the notion of Neural Controlled Differential Equations  [29].
[…]
Figure 1: Schematic of the proposed DAE Timestepper.
Training the algebraic update sub-task is accomplished by formulating a PINN-style penalty loss [5] in which known constraints are used to ‘penalize’ the discrepancy between constraint satisfaction and predicted states. These penalties are assessed in a multi-objective optimization loss function:
$\mathcal{L}(\Theta)=\lambda_{1}\mathcal{L}_{\text{residual}}(X,\hat{X};\Theta)+\lambda_{2}\mathcal{L}_{\text{constraints}}(X,\hat{X};\Theta),$ (10)
where $\mathcal{L}_{\text{residual}}$ is a measure of the residual between the training dataset and the model’s output, $\mathcal{L}_{\text{constraints}}$ is a regularization term penalizing the constraint violations. The $\lambda$’s represent weightings for balancing the multi-objective loss function. We optimize model parameters $\Theta$ with an off-the-shelf stochastic gradient-based optimizer [30] to solve $\Theta^{*}=\underset{\Theta}{\text{argmin}}~{}\mathcal{L}(\Theta)$.
For this work, we choose minimal architectures for ease of implementation and exposition, as outlined in Alg. 1. We use a first-order explicit timestepper for the ODE integration and use a DNN as a neural surrogate of an algebra solver.
Data: $\hat{X}=\{(\hat{x}^{(0)},\hat{y}^{(0)}),...,(\hat{x}^{(N)},\hat{y}^{(N)})\}$: Time series dataset with spacing $\Delta t$
Parameters : $\theta_{f},\theta_{h}$: Tunable parameters for mappings $f$ and $h$
/* Pick appropriate integrator */
1 Function ODESolve($f,\{x,y,u\}$):
2       $x^{*}\longleftarrow x+\Delta t\times f\left(x,y,u;\theta_{f}\right)$
3       return $x^{*}$
/* Sequential sub-tasks */
4 Function DAESolve($f,h,\{x,y,u\}$):
5       $y^{*}\longleftarrow h\left(x,y,u;\theta_{h}\right)$
6       $x^{*}\longleftarrow\textnormal{{ODESolve}}\left(f,\{x,y^{*},u\};\theta_{f}\right)$
7       return $(x^{*},y^{*})$
/* Training loop */
8 while not converged do
       /* Forward pass */
9       $X\longleftarrow\{(\hat{x}^{(0)},\hat{y}^{(0)})\}$
10       for k $\in 1,\dots,N$ do
11             $X\longleftarrow$Concat ($X$,
12             DAESolve($f,h,\{x^{(k-1)},y^{(k-1)},u^{(k-1)}\}$) )
13
      /* Evaluate loss */
14       $\mathcal{L}\longleftarrow$ Criterion ( $X,\hat{X}$ )
15       if converged then
16             break;
17
18      else
19             obtain gradients;
20             update model weights $\{\theta_{f},\theta_{h}\}$;
21
22
Algorithm 1 Neural DAE training
IV NUMERICAL CASE STUDIES
All experiments were performed using the NeuroMANCER Scientific Machine Learning (SciML) package [31] built on top of PyTorch [32] [^1] Repository containing all training and visualization scripts will be open-sourced and publicly available by the time of publication.. Each presented model was trained using an Intel I9-9980 CPU laptop with a wall-clock training time on the order of 5 minutes.
IV-A Tank-Manifold Property Inference
IV-A1 Problem Description
A manifold is a physical realization of a conservation relationship that manifests as an algebraic constraint in a DAE. Figure  2 depicts a pair of tanks fed by a manifold. A pseudo-infinite reservoir supplies the pump. At the manifold, the inflow is exactly equal to the sum of the outflows; that is, $y_{\text{in}}=y_{1}+y_{2}$, where $y$ represents volumetric flow rates and the subscripts $1$ and $2$ denote the two exit streams of the manifold. In this example, the two tanks have different area-height profiles but have a common datum such that the column heights are identical. As a fluid is pumped into the two tanks, the flow rates in the two streams varies to ensure (i) the conservation relationship of the manifold and (ii) that each of the tanks have identical pressure head, i.e $x_{1}=x_{2}$, where $x$ is the height of the fluid in the respective tank. The full DAE that describes the evolution of this system is given as:
$\begin{split}\frac{dx_{1}}{dt}&=\frac{y_{1}}{\phi_{1}(x_{1})}\\ \frac{dx_{2}}{dt}&=\frac{y_{2}}{\phi_{2}(x_{2})}\end{split}$ (11)
$0=y_{\text{in}}-y_{1}-y_{2}\quad$ $0=x_{1}-x_{2},$
where $\phi_{1}$ and $\phi_{2}$ denote the area-height profiles for the two tanks. The solution of Eq.  11 can be obtained with many off-the-shelf standard scientific computing packages. For this study, the area-height profiles are set to $\phi_{1}(x)=3.0$ and $\phi_{2}(x)=\sqrt{x}+0.1$ respectively. The goal of the learning task is to recover $\phi_{2}(x)$ given observations of tank heights $x$ and pump volumetric flow rates $y$. Note that if the task were to recover both $\phi_{1}(x)$ and $\phi_{2}(x)$ simultaneously, the problem would be ill-posed in the sense that there does not exist a unique set of tank profiles that recovers the behavior seen in observations.
[…]
Figure 2: Tank-manifold-pump schematic. A reservoir supplies fluid to a pump-manifold system. The tanks have different area-height profiles but a common datum. Conservation is enforced at the manifold.
IV-A2 Model Construction
In the system described by Eq.  11, the differential states are column heights $x_{1}$ and $x_{2}$ and the algebraic states are the volumetric flows $y_{1}$ and $y_{2}$. The input flow is treated as an exogenous control input, i.e., $u=y_{\text{in}}$.
We aim to approximate the area-height profile of Tank #1 as well as construct a surrogate for the behavior of the manifold such that (i) fluid is conserved and (ii) the fluid column height constraint is satisfied. There are two neural networks that we seek to train: $NN_{1}:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$, an approximation of the sought area-height profile $\phi_{1}$, and $NN_{2}:\mathbb{R}^{4}\rightarrow(0,1)$, the surrogate of the behavior of the manifold. Each are implemented as a feed-forward neural network with one hidden layer of 5 nodes, each with sigmoid activation functions. The output layer of $NN_{1}$ is linear while the output layer of $NN_{2}$ is sigmoidal, constraining the output to $(0,1)$.
IV-A3 Optimization Problem
A time series dataset is constructed by simulation of the system given in Eq.  11 regularly sampled on the interval $t\in[0,500]$ with $\Delta t=1.0$ and with a constant inflow of $y_{\text{in}}=u=0.5$:
$\hat{X}=\{(\hat{x}^{(0)},\hat{y}^{(0)}),...,(\hat{x}^{(N)},\hat{y}^{(N)})\},$ (12)
where $N$ is the time series prediction horizon and the $\widehat{\cdot}$ denotes membership to the dataset. The optimization problem seeks the minimum of the differential state trajectory reconstruction loss and algebraic state reconstruction loss when predicted by our $\operatorname{DAESolve}$:
$\begin{split}\underset{\theta_{1},\theta_{2}}{\text{min}}&~{}~{}\lambda_{1}\sum_{k=1}^{N}||x^{(k)}-\hat{x}^{(k)}||_{2}^{2}+\lambda_{2}\sum_{k=1}^{N}||y^{(k)}-\hat{y}^{(k)}||_{2}^{2}\\ \text{s.t.~{}~{}}&\{x^{(k+1)},y^{(k+1)}\}\\ ~{}&=\operatorname{DAESolve}\left(f,h,\{x^{(k)},y^{(k)},u^{(k)}\};\{\theta_{1},\theta_{2}\}\right),\\ \end{split}$ (13)
The functions $f$ and $h$ are defined as:
$\begin{split}f&:\begin{cases}\frac{y_{1}}{3}\\ \frac{y_{2}}{\operatorname{NN}_{1}(x_{2};\theta_{1})},\\ \end{cases}\\ h&:\begin{cases}u\operatorname{NN}_{2}(x_{1},x_{2},y_{1},y_{2};\theta_{2})\\ u\left(1-\operatorname{NN}_{2}(x_{1},x_{2},y_{1},y_{2};\theta_{2})\right),\\ \end{cases}\\ \end{split}$ (14)
Note that the mapping $h$ defines a convex combination at the manifold: the outflows $y_{1}$ and $y_{2}$ will sum to $u$ by construction.
For all examples contained in this work, the optimizer is Adam [30] with fixed learning rate of 0.001. Training is deemed ‘converged’ after 20,000 epochs or until an early stop is triggered after 20 epochs of increasing loss.
IV-A4 Results
The time series data for the tank heights and volumetric flows are given in Figs. 3(a) and 3(b) respectively. The trained model is able to accurately reproduce the time series data seen in training. Additionally, it is able to accurately learn the model of the unknown area-height profile of Tank #2. Figure 4(a) depicts the ground-truth area-height profile and the model’s reconstruction. Furthermore, to demonstrate the generalization, the trained model is tested on an unseen inlet mass flow rate; $u(t)=\frac{1}{2}+\frac{1}{4}\sin{\left(\frac{t}{100}\right)}$. The resulting time series of volumetric flows from the trained model rollout is compared with the ground-truth conditions in Fig. 4(b). Table-I gives the Mean Squared Error (MSE) of the trained neural DAE model.
Figure 3: (a) Tank heights. (b) Volumetric flow.
Figure 4: Trained model (a) Area-height relationships (b) Volumetric flow time series.
Extrapolation to a new flow condition demonstrates the ability to (i) endogenize the behaviors of the components in the network and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the network.
TABLE I: Mean Square error for the differential state and algebraic states for tanks-manifold-pump experiment.
Manifold
Mean Square Error
Tanks Height 9e-03
Volumetric Flow 2e-02
Area-Height 6e-03
Unseen inlet mass-flow 1e-01
IV-B Tank Network Modeling
IV-B1 Problem Description
Figure 5 expands upon the tank-manifold problem of Fig. 2 by adding additional tanks, feedback mechanisms, and ‘closing the loop’ in the network flow. In this example, the pump flow rate depends on the fluid heights of Tank #1 and the reservoir. Tanks #1 and #2 have a height constraint (equivalent pressure head) and are fed by a common manifold. The outlets of tanks #1 and #3 depend on the square root of the fluid column height in the respective tanks. The full system can be described by the DAE:
$\begin{split}\frac{dx_{1}}{dt}&=\frac{1}{\phi_{1}(x_{1})}\left(y_{1}-y_{3}\right)\\ \frac{dx_{2}}{dt}&=\frac{1}{\phi_{2}(x_{2})}\left(y_{2}\right)\\ \frac{dx_{3}}{dt}&=\frac{1}{\phi_{3}(x_{3})}\left(y_{3}-y_{4}\right)\\ \frac{dx_{4}}{dt}&=\frac{1}{\phi_{4}(x_{4})}\left(y_{4}-y_{0}\right)\\ \end{split}$ (15)
$0=y_{0}-p(x_{1},x_{4})\quad$ $0=y_{0}-y_{1}-y_{2}$
$0=y_{3}-\alpha_{1}\sqrt{x_{1}}\quad$ $0=y_{4}-\alpha_{2}\sqrt{x_{3}}$
$0=x_{1}-x_{2}\quad$
where $x$ denotes the fluid column height for tanks #1 through #4 (the reservoir), $p$ is a pump response function, $y$ represents volumetric flows, and $\alpha$ describe the discharge coefficients of the tank outlets. The differential states are $x=\{x_{1},x_{2},x_{3},x_{4}\}\in\mathbb{R}^{4}$ and the algebraic states are $y=\{y_{0},y_{1},y_{2},y_{3},y_{4}\}\in\mathbb{R}^{5}$.
In this example, several characteristics make the dynamics more complex. First, the system is closed, as there is no material transfer into or out of the system. Second, there is the potential for backflow through the manifold from Tank #2. Lastly, the feedback mechanism from the tank level transducers to the pump controller is such that oscillations can occur.
[…]
Figure 5: Networked system defined by tanks, pipes, a manifold, and a pump. The system exhibits oscillatory dynamics for some parameter choices and sensing schemes.
IV-B2 Optimization Problem
The system given in Eq. (15) was simulated with the following specifications:
$\phi_{1}(x_{1})=2.0\quad$ $\phi_{2}(x_{2})=1.0$
$\phi_{3}(x_{3})=1.0\quad$ $\phi_{4}(x_{4})=10.0$
$p(x_{1},x_{4})=0.1x_{1}x_{4}\quad$ $\alpha_{1},\alpha_{2}=0.1$ (16)
A single-trajectory dataset $\hat{X}$ was constructed through sampling this simulation at $\Delta t=0.1$ for $t\in[0,20]$. The optimization problem definition is identical to that of Eq. 13. For this problem, it is reasonable to assume that the basic physics for a gravity-fed pipe network are known, which provides a starting point for constructing a model. However, discharge coefficients and quantifying feedback mechanisms are not as straightforward. Here, we attempt to approximate (i) the Level Controller (LC) dynamics for pump control with a neural network $\operatorname{NN}_{1}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{1}$, (ii) manifold dynamics with with a second neural network $\operatorname{NN}_{2}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{1}$, and (iii) the discharge coefficients $\alpha_{1}$ and $\alpha_{2}$. The mappings $f$ and $h$ are defined as:
$\begin{split}f&:\begin{cases}\frac{y_{1}-y_{3}}{\phi_{1}(x_{1})}\\ \frac{y_{2}}{\phi_{2}(x_{2})}\\ \frac{y_{3}-y_{4}}{\phi_{3}(x_{3})}\\ \frac{y_{4}-y_{0}}{\phi_{4}(x_{4})},\\ \end{cases}\\ h&:\begin{cases}\operatorname{NN}_{1}(x_{1},x_{4};\theta_{1})\\ y_{0}\operatorname{NN}_{2}(x_{2},x_{3};\theta_{2})\\ y_{0}(1-\operatorname{NN}_{2}(x_{2},x_{3};\theta_{2}))\\ \alpha_{1}\sqrt{x_{1}}\\ \alpha_{2}\sqrt{x_{3}}.\\ \end{cases}\end{split}$ (17)
For this example, both $\operatorname{NN}_{1}$ and $\operatorname{NN}_{2}$ are simple multi-layer perceptrons each with a two hidden layers of size 30 and ReLU activations.
IV-B3 Results
Figure 6(a) details the time history of the four tank heights for both the ground truth data as well as the open-loop model rollout. At these process conditions, the pump delivers an oscillatory flow to the manifold, which our model is able to capture. Additionally, the tuned model is able to infer the values for the discharge coefficients: for the case of zero noise the recovered values are $\alpha_{1}=0.1027$ and $\alpha_{2}=0.1024$.
To demonstrate generalizability, the trained model is deployed on an unseen area-height tank profile; for this, we redefine $\phi_{1}(x_{1})=1.0$. In doing so, the system’s limit cycle is destroyed, and the true dynamics approach a steady state. Figure  6(b) demonstrates this extrapolation for the same time horizon. The reconstructed time series data matches the behavior of the unseen data: there exist damped oscillations that decay to a steady state value. Table-II shows the MSE for the neural DAE models trained for the tank network problem.
Noise tolerance is an important consideration in the evaluation of system identification tasks, especially in settings with nonlinearities and feedback mechanisms. For this example, we investigate noise tolerance by introducing Gaussian white noise. Figures  7(a) and  7(b) show similar training and extrapolation tasks for the Signal-to-Noise Ratio (SNR) of $20$ dB. Despite the significant noise, the learned dynamics are faithful to the underlying physics, as demonstrated in the extrapolation task, and result in an MSE of $027e-$ 02.
Figure 6: Tank Network problem : (a) Ground truth data vs Neural-DAE model prediction. (b) Extrapolation case: inferring behavior for unseen component parameters using Neural-DAE.
Figure 7: Tank Network problem : (a) noisy data vs Neural-DAE model prediction. (b) Extrapolation case: inferring behavior for unseen component parameters using Neural-DAE.
TABLE II: Mean Square error for the differential state and algebraic states for tanks network experiment.
Trained Model Type Tank-Network Experiment
Mean Square Error
No Noise Tanks Height 1e-02
Unseen initial Data : Tanks Height 6e-02
20 dB Noise Tanks Height 1.03
Unseen initial Data : Tanks Height 7e-2
V CONCLUSIONS AND FUTURE WORKS
This work presents a framework for the data-driven modeling of Differential Algebraic Equations (DAEs). Specifically, we introduce a novel method called Neural Differential Algebraic Equations (NDAEs) that extends the notion of Universal Differential Equations to endogenize the evolution of algebraic relationships. Inspired by Lie-Trotter operator splitting, we achieve this by splitting the integration of system states into two sequential sub-tasks: first, updating algebraic states via a differentiable surrogate of an algebra solver, followed by a conventional temporal integration step for a system of Ordinary Differential Equations.
The effectiveness of these methods was demonstrated by (i) modeling networked dynamical systems with underlying constraints and conservation relationships and (ii) showing generalization to unseen process conditions or component variation. We demonstrate that the proposed NDAE method is robust to noise and possesses the extrapolation ability to (i) learn the behaviors of the DAE system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the DAE system.
While the methods and example application problems were successful in this work, future research and development are needed in several areas. First, a central assumption of our method is that the underlying dynamics of the systems of interest are sufficiently smooth, i.e., not containing jump discontinuities or shocks, such that the underlying dynamics are non-stiff. As presented, our method leverages an explicit fractional-step approach to decoupling the evolution of the differential and algebraic states. To increase the ability of the method to handle stiff dynamics or discontinuities, moving to an implicit integration scheme may be warranted. Second, further analysis is needed to investigate the interplay between the modeling task and the index of the underlying DAE (for example, higher-order DAEs that are characteristic of design or parameter optimization problems). Lastly, we have made no claims about computational scalability nor partial observability, both of which are critical for modeling real-world complex systems with hundreds of degrees of freedom. We reserve these topics for future work.
ACKNOWLEDGMENTS
This research was supported by the U.S. Department of Energy through the Building Technologies Office under the Advancing Market-Ready Building Energy Management by Cost-Effective Differentiable Predictive Control projects. PNNL is a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL0-1830.